OpenStudy (anonymous):

solve for tan theta = 2sin theta

OpenStudy (anonymous):

i got that far allready, but the answer in the back of the book is telling me 0, Pi, Pi/3, -Pi/3.

OpenStudy (mimi_x3):

can you tell me what you did?

OpenStudy (anonymous):

i have'nt done anything, only what you have done, i need to know how to continue the equation and find those values, i guess using the circle quadrant or something like that

OpenStudy (unklerhaukus):

2sinθ=2sinθcosθ ???

OpenStudy (mimi_x3):

wooopppss..i looked at it wrongly; i thought it sin2x sorry

OpenStudy (anonymous):

\[\tan(\theta) = 2\sin(\theta) \implies \frac{\sin(\theta)}{\cos(\theta)} = 2 \sin(\theta) \implies 2\cos(\theta) = 1\]

OpenStudy (mimi_x3):

then \[ tan\theta = 2sin\theta => tan\theta - 2sin\theta\] \[\frac{\sin\theta}{\cos\theta} -2\sin\theta = 0 => \frac{\sin\theta}{\cos\theta} - \frac{2\sin\theta*\cos\theta}{1*\cos\theta} = 0 \frac{\sin\theta-\sin2\theta}{\cos\theta} =0\] lol, i went the long way woops.

OpenStudy (anonymous):

\[\cos(\theta) = \frac{1}{2} \implies \cos(\theta ) = \cos(\frac{\pi}{3}) \implies \theta = 2 n \pi \pm \frac{\pi}{3}\]

OpenStudy (anonymous):

tanθ=2sinθ sinθ/cosθ=2sinθ sinθ=2sinθcosθ 2sinθcosθ-sinθ=0 sinθ(2cosθ-1)=0 sinθ=0 or cosθ=1/2 so, θ=0,pi and θ=pi/3 or -pi/3

OpenStudy (anonymous):

Where \(n \in \mathbb{Z}\).. Once can find all the solution by putting n = -2, -1 0, 1,2 etc etc. I mean to say that integer values..

OpenStudy (anonymous):

thanks very much that's right but how do i show that on the unit circle

OpenStudy (anonymous):

abhiphardhu

OpenStudy (anonymous):

@abhipardhu